OpenStudy (anonymous):

can someone Simplify (2-sqroot8) (3+sqroot2)

OpenStudy (anonymous):

(2-2√(2)) (3+√(2))

OpenStudy (anonymous):

(2 - √8) (3 + √2) = (2 - 2√2) (3 + √2) = 6 + 2√2 -6√2 -4 = 2 - 4√2

OpenStudy (anonymous):

thank you very much